# ReasoningV: Efficient Verilog Code Generation with Adaptive Hybrid Reasoning Model

### v1

@@ Abstract
Large Language Models (LLMs) have advanced Verilog code generation significantly, yet face challenges in data quality, reasoning capabilities, and computational efficiency. This paper presents ReasoningV, a novel model employing a hybrid reasoning strategy that integrates trained intrinsic capabilities with dynamic inference adaptation for Verilog code generation. Our framework introduces three complementary innovations: (1) ReasoningV-5K, a high-quality dataset of 5,000 functionally verified instances with reasoning paths created through multi-dimensional filtering of PyraNet samples; (2) a two-stage training approach combining parameter-efficient fine-tuning for foundational knowledge with full-parameter optimization for enhanced reasoning; and (3) an adaptive reasoning mechanism that dynamically adjusts reasoning depth based on problem complexity, reducing token consumption by up to 75% while preserving performance. Experimental results demonstrate ReasoningV’s effectiveness with a pass@1 accuracy of 57.8% on VerilogEval-human, achieving performance competitive with leading commercial models like Gemini-2.0-flash (59.5%) and exceeding the previous best open-source model by 10.4 percentage points. ReasoningV offers a more reliable and accessible pathway for advancing AI-driven hardware design automation, with our model, data, and code available at https://github.com/BUAA-CLab/ReasoningV.
@@ I Introduction
    ReasoningV-5K Dataset: A high-quality, functionally verified dataset of 5,000 problem-solution-test triplets derived from 690K samples, tackling the data quality deficiency.
@@
Experiments show ReasoningV achieves state-of-the-art (SOTA) results among open-source models, notably reaching 57.8% pass@1 on VerilogEval-Human[11], surpassing the previous best by 10.4 percentage points [15]. This work demonstrates that targeted methodological ingenuity in data curation, training, and inference can markedly improve the fidelity and practicality of AI-generated hardware descriptions, advancing reliable and resource-efficient AI tools for the hardware development lifecycle. The remainder of this paper details our methodology, results, and related work.
@@ II-A LLMs and Verilog Generation Challenges
PyraNet GitHub+LLM 690K+ Syntax-checked, quality-ranked
MG-Verilog Open-Source 11K+ Multi-grain descriptions
RTLCoder LLM-gen 80K+ ABV/FPV subset available
CodeV GitHub 165K Multi-level summaries
VerilogEval HDLBits 330 Testbench evaluation
@@ II-B Verilog Datasets and the Need for Quality
ReasoningV-5K addresses this gap. Derived from filtering PyraNet, it provides 5,000 high-quality “problem-solution-test” triplets. Crucially, every solution is functionally verified via simulation using its corresponding testbench, and each sample includes a distilled reasoning path. This combination directly tackles data quality, verification, and reasoning deficits needed for more reliable Verilog generation models.
@@ II-C Verilog Generation Models and Reasoning Limitations
ReasoningV adopts a distinct hybrid reasoning strategy. Unlike purely prompt-driven methods, it invests in training intrinsic reasoning using the ReasoningV-5K dataset and a two-stage training method. This learned capability is managed by an adaptive inference mechanism (Sec III-C) for efficiency. This aims for greater robustness than prompt-elicited reasoning and more targeted enhancement than standard tuning or RL focused mainly on code structure or basic correctness.
@@ III Methodology
This section outlines the ReasoningV framework for Verilog code generation. As shown in Figure 1, our approach comprises three core components: constructing the ReasoningV-5K dataset, implementing a two-stage training method, and developing an adaptive reasoning mechanism. These elements systematically address poor data quality, limited reasoning capabilities, and computational inefficiency through data refinement, staged training, and dynamic reasoning optimization.
Figure 1: Overall architecture of the ReasoningV framework.
@@ III-A High-Quality Reasoning Dataset Construction
High-quality training data is vital for effective HDL generation. We constructed ReasoningV-5K, a rigorously verified Verilog reasoning dataset containing 5,000 high-quality samples, through multi-stage filtering and verification processes designed to enhance data reliability and enforce design best practices. As illustrated in Figure 2, our construction process consists of three main steps: (1) Multi-Dimensional Filtering of PyraNet samples, (2) Reasoning Enhancement with DeepSeek-R1, and (3) Final Dataset Assembly with functional testbench verification.
@@ III-A3 Dataset Verification and Structure
We adopted a “problem-description-reasoning_path-solution-testbench” format. Each entry includes the problem description, a detailed reasoning path (distilled using DeepSeek-R1, see Figure 3), the Verilog solution, and a corresponding functional testbench (also generated with DeepSeek-R1 guidance).
@@
Crucially, functional verification was performed by simulating each solution against its testbench using Icarus Verilog [33]. Samples failing simulation (approx. 86% of the 12K candidates) were discarded. This prioritization of functional accuracy resulted in the final ReasoningV-5K dataset: 5,000 high-quality, functionally verified instances with reasoning paths, exhibiting high information density (CR 4.56).
We used the OriGen dataset [15] (220K) for initial commonsense training (Stage 1) and reserved ReasoningV-5K for reasoning enhancement (Stage 2). Our filtering reduced the initial 690K samples by approx. 99.3%, significantly improving quality and ensuring functional correctness per sample.
@@ III-C1 Judge Adapter Training
We systematically sampled 10,000 representative questions from the training corpus and constructed a specialized difficulty-classification dataset through knowledge distillation from the DeepSeek-V3 model. This engineered dataset was then employed to perform LoRA training on our base model, yielding a Judge Adapter specifically optimized for hierarchical question difficulty assessment (categorizing into “Easy”, “Medium”, or “Hard”) through parameter-efficient fine-tuning.
III-C2 Reasoning Mode Selection and Application Logic
As illustrated in Figure 1 (Innovation 3), the framework uses the Judge Adapter’s output to dynamically select reasoning modes and allocate resources:
Operational Phases and Task Division
The system operates in two modes:
  1. 1.
    Difficulty Assessment: Activate Judge Adapter. Input: Problem description. Output: Difficulty category (“Easy”/“Medium”/“Hard”). Minimal overhead ($<$1% tokens). Purpose: Quick complexity estimation.
  2. 2.
    Code Generation: Disable Judge Adapter. Use full ReasoningV model. Input: Problem description (+ optional reasoning prompt). Output: Reasoning path (if applicable) and Verilog code. Purpose: Generate solution based on selected strategy.
Dynamic Parameter Adjustment and Mode Selection
Based on assessed difficulty, one mode is selected:
@@ III-C2 Reasoning Mode Selection and Application Logic
Resource Optimization
This adaptive approach significantly optimizes resource use. The lightweight Judge Adapter is active only briefly. The conditional generation strategy allocates compute efficiently—more for hard problems, less for easy ones—optimizing throughput and cost compared to fixed-depth reasoning approaches.
@@ IV-B Main Results
Verilog-Specific Models RTLCoder-DeepSeek 41.6 50.1 53.4 61.2 76.5 81.8 - -
@@ IV-C2 Analysis of Adaptive Reasoning Mechanism
Figure 4 illustrates the relationship between model size (parameters) and performance (pass@1 on VerilogEval-Human) for the base Qwen2.5-Coder models compared to our ReasoningV variants.

### v3

@@ Abstract
Large Language Models (LLMs) have advanced Verilog code generation significantly, yet face challenges in data quality, reasoning capabilities, and computational efficiency. This paper presents ReasoningV, a novel model employing a hybrid reasoning strategy that integrates trained intrinsic capabilities with dynamic inference adaptation for Verilog code generation. Our framework introduces three complementary innovations: (1) ReasoningV-5K, a high-quality dataset of 5,334 functionally verified instances with reasoning paths created through multi-dimensional filtering of PyraNet samples; (2) a two-stage training approach combining parameter-efficient fine-tuning for foundational knowledge with full-parameter optimization for enhanced reasoning; and (3) an adaptive reasoning mechanism that dynamically adjusts reasoning depth based on problem complexity, reducing token consumption by up to 75% while preserving performance. Experimental results demonstrate ReasoningV’s effectiveness with a pass@1 accuracy of 57.8% on VerilogEval-human, achieving performance competitive with leading commercial models like Gemini-2.0-flash (59.5%) and exceeding the previous best open-source model by 10.4 percentage points. ReasoningV offers a more reliable and accessible pathway for advancing AI-driven hardware design automation, with our model, data, and code available at https://github.com/BUAA-CLab/ReasoningV.
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication.
Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
    ReasoningV-5K Dataset: A high-quality, functionally verified dataset of 5,334 problem-solution-test triplets derived from 690K samples, tackling the data quality deficiency.
@@
Experiments show ReasoningV achieves state-of-the-art (SOTA) results among open-source models, notably reaching 57.8% pass@1 on VerilogEval-Human[11], surpassing the previous best by 10.4 percentage points [15]. This work demonstrates that targeted methodological ingenuity in data curation, training, and inference can markedly improve the fidelity and practicality of AI-generated hardware descriptions, advancing reliable and resource-efficient AI tools for the hardware development lifecycle. The remainder of this paper details our related work, methodology, and results.
@@ II-A LLMs and Verilog Generation Challenges
PyraNet[16] GitHub+LLM 690K+ Syntax-checked, quality-ranked
MG-Verilog[17] Open-Source 11K+ Multi-grain descriptions
RTLCoder[10] LLM-gen 80K+ ABV/FPV subset available
CodeV[18] GitHub 165K Multi-level summaries
VerilogEval[11] HDLBits 330 Testbench evaluation
@@ II-B Verilog Datasets and the Need for Quality
ReasoningV-5K addresses this gap. Derived from filtering PyraNet[16], it provides 5,334 high-quality “problem-solution-test” triplets. Crucially, every solution is functionally verified via simulation using its corresponding testbench, and each sample includes a distilled reasoning path. This combination directly tackles data quality, verification, and reasoning deficits needed for more reliable Verilog generation models.
@@ II-C Verilog Generation Models and Reasoning Limitations
ReasoningV adopts a distinct hybrid reasoning strategy. Unlike purely prompt-driven methods, it invests in training intrinsic reasoning using the ReasoningV-5K dataset and a two-stage training method. It aims for greater robustness than prompt-elicited reasoning and more targeted enhancement than standard tuning or RL focused mainly on code structure or basic correctness.
@@ III Methodology
This section outlines the ReasoningV framework for Verilog code generation. As shown in Fig.1, our approach comprises three core components: constructing the ReasoningV-5K dataset, implementing a two-stage training method, and developing an adaptive reasoning mechanism. These elements systematically address poor data quality, limited reasoning capabilities, and computational inefficiency through data refinement, staged training, and dynamic reasoning optimization.
@@ III-A High-Quality Reasoning Dataset Construction
High-quality training data is vital for effective HDL generation. We constructed ReasoningV-5K, a rigorously verified Verilog reasoning dataset containing 5,334 high-quality samples, through multi-stage filtering and verification processes designed to enhance data reliability and enforce design best practices. As illustrated in Fig. 2, our construction process consists of three main steps: (1) Multi-Dimensional Filtering of PyraNet samples[16], (2) Reasoning Enhancement with DeepSeek-R1[32], and (3) Final Dataset Assembly with functional testbench verification.
@@ III-A3 Dataset Verification and Structure
We adopted a “problem-description-reasoning_path-solution-testbench” format. Each entry includes the problem description, a detailed reasoning path (distilled using DeepSeek-R1, see Fig 3), the Verilog solution, and a corresponding functional testbench (also generated with DeepSeek-R1 guidance).
@@
Crucially, functional verification was performed by simulating each solution against its testbench using Icarus Verilog [34]. Samples failing simulation (approx. 23.3% of the 7K candidates) were discarded. This prioritization of functional accuracy resulted in the final ReasoningV-5K dataset: 5,334 high-quality, functionally verified instances with reasoning paths, exhibiting high information density (CR 4.56).
Our filtering reduced the initial 690K samples by approx. 99.3%, significantly improving quality and ensuring functional correctness per sample.
@@ III-C1 Judge Adapter Training
We systematically sampled 10,000 representative questions from the training corpus and constructed a specialized difficulty-classification dataset through knowledge distillation from the DeepSeek-V3 model. This engineered dataset was then employed to perform LoRA training on our base model, yielding a Judge Adapter specifically optimized for hierarchical question difficulty assessment. The adapter categorizes each input into one of three difficulty levels: “Easy”, “Medium”, or “Hard”, enabling fine-grained control over downstream reasoning strategies.
To better align inference strategies with difficulty levels, we introduced dynamic mode selection based on the Judge Adapter’s assessment:
@@
III-C2 Reasoning Mode Selection and Application Logic
As illustrated in Fig.1 (Innovation 3), the framework uses the Judge Adapter’s output to dynamically select reasoning modes and allocate resources:
Operational Phases and Task Division The system operates in two modes:
  • Code Generation: Disable Judge Adapter. Use full ReasoningV model.
    • Input: Problem description (+ optional reasoning prompt).
    • Output: Reasoning path (if applicable) and Verilog code.
    • Purpose: Generate solution based on selected strategy.
Resource Optimization This adaptive approach significantly optimizes resource use:
  • The lightweight Judge Adapter is active only briefly
  • Conditional generation allocates compute efficiently:
    • More resources for hard problems
    • Fewer resources for easy ones
  • Optimizes throughput and cost compared to fixed-depth reasoning approaches
@@ IV-B Main Results
Verilog-Specific Models RTLCoder-DeepSeek 41.6 50.1 53.4 61.2 76.5 81.8 33.5 43.9
@@ IV-C2 Analysis of Adaptive Reasoning Mechanism
Fig. 4 illustrates the relationship between model size (parameters) and performance (pass@1 on VerilogEval-Human) for the base Qwen2.5-Coder models compared to our ReasoningV variants.